OpenStudy (wonderwomen):

what is the missing fraction? 1/6+?=11/18 A) 10/6 B) 4/9 C) 2/3 D) 10/18

OpenStudy (anonymous):

Any idea about algebra!

OpenStudy (wonderwomen):

do not be rude!

OpenStudy (anonymous):

Rude???

OpenStudy (wonderwomen):

whatever lets just get back to the question

OpenStudy (anonymous):

Did u see what macc... asked!!

OpenStudy (wonderwomen):

maccrea?

OpenStudy (anonymous):

Well as for ur question lets suppose the unknown is x, so the equation becomes '1/6+x= 11/18'. Now 'x= 11/18 - 1/6' Hope u can do it from here :)

OpenStudy (wonderwomen):

it is B) 4/9 ?

OpenStudy (anonymous):

Yeah! u guys doing algebra in school

OpenStudy (wonderwomen):

thanks